OpenStudy (anonymous):

whats a line parallel to y=2x-9 whats a line perpendicular to y=2x-9

OpenStudy (anonymous):

parallel- y=2x-6 and perpendicular- y=1/2x-9

geerky42 (geerky42):

y=-1/2x-9*

OpenStudy (anonymous):

thanks I forgot that sneaky negative

OpenStudy (anonymous):

can I see the work for It so I can understand it

geerky42 (geerky42):

This may help you: http://www.mathsisfun.com/algebra/line-parallel-perpendicular.html

OpenStudy (anonymous):

for making a line parallel you keep the same slope and just change the y-intercept. for making a line perpendicular you take the inverse of the slope and make it negative

geerky42 (geerky42):